A genetic experiment involving peas yielded one sample of offspring consisting of 446446 green peas and 155155 yellow peas. Use a 0.010.01 significance level to test the claim that under the same​ circumstances, 2424​% of offspring peas will be yellow. Identify the null​ hypothesis, alternative​ hypothesis, test​ statistic, P-value, conclusion about the null​ hypothesis, and final conclusion that addresses the original claim. Use the​ P-value method and the normal distribution as an approximation to the binomial distribution.
Answer #1

H0: Null Hypothesis: P = 0.24

HA: Alternative Hypothesis: P eq 0.24

n = Sample Size = 446 + 155 = 601

p = Sample Proportion = 155/601 = 0.2579

P = Population proportion = 0.24

Q = 1- P = 0.76

SE = sqrt{PQ/n}

V0.24 × 0,76/601 0.0174

Test statistic is:
Z = (p - P)/SE

= (0.2579 - 0.24)/0.0174 = 1.0287

Table of Area Under Standard Normal Curve gives area = 0.3485

So,

P - Value = (0.5 - 0.3485) X 2 = 0.3030

alpha = 0.01

Since P - Value = 0.3030 is greater than alpha = 0.01, the difference is not significant. Fail to reject null hypothesis.

Conclusion:

The data support the claim that under the same circumstances, 24% of offspring peas will be yellow.